OpenStudy (anonymous):

ind the x- and y-intercepts for the equation. Please show all of your work. y = -2x – 8

OpenStudy (anonymous):

for y-intercept, put x=0 you get y = -8 so y-intercept is -8 for x -intercept, put y =0 0 = -2x -8 8 = -2x 8/-2 = x -4 = x so x = -4 is the x intercept